# Impartial Games: a Challenge for Reinforcement Learning

### v5

@@ 6 Concluding Remarks and Conjectures
From a cognitive science perspective, the difference between games like Chess/Go and Nim is not merely one of complexity but of the very nature of optimal strategy. Mastering Chess and Go primarily involves associative pattern matching, heuristic evaluation of board configurations, and deep subsymbolic calculation. A human Grandmaster’s mastery is built upon millions of learned patterns, making it an ideal domain for reinforcement learning agents that excel at function approximation on large, but structured, data distributions. In contrast, Nim is governed by a single, elegant mathematical principle: the nim-sum. A human’s mastery of Nim is non-incremental; it is achieved not through extensive pattern recognition but through a single, abstract insight into combinatorial game theory. This disparity highlights a crucial distinction: AlphaZero excels at learning through association, but fails when a problem requires a form of symbolic reasoning that cannot be implicitly learned from the correlation between game states and outcomes.
@@
Our work empirically demonstrates that this theoretical vulnerability becomes a tangible, catastrophic failure mode in practice. While the problem of a neural network’s difficulty in learning parity from a uniform distribution is a well-established theoretical finding, our contribution is to show how this representational bottleneck manifests and breaks the AlphaZero learning loop in a real-world, self-play setting. The motivation behind our research is not to simply identify a known weakness, but to meticulously dissect its consequences and chart a path for extending the AlphaZero paradigm. By analyzing the breakdown of the policy and value networks on specific Nim positions, we have revealed how a critical representational gap can trigger a cascade of failures.
@@
The core of the problem, as our analysis shows, is a breakdown in the crucial positive feedback loop that underpins reinforcement learning. The agent’s performance on small boards, while seemingly successful, is based on a precarious "champion" notion of mastery—memorizing a few optimal opening moves. This allows the agent to steer the game into a sub-space it has seen frequently. However, as the number of heaps increases to just seven, the policy network’s predictions become near-uniform, and the value network’s assessments drift to a coin-toss probability. This lack of a strong initial policy prevents the MCTS from exploring the state space effectively, leaving it unable to find winning moves, even with millions of simulations. This is not a failure of calculation; it is a failure of learned guidance. Crucially, this happens even on boards where the state space is still small enough that many states would be revisited during training. In this scenario, the learning process becomes completely derailed, as the agent fails to generalize to any new positions, regardless of whether the parity function is complex. When the state space becomes truly vast (e.g., with 50+ heaps), the difficulty of learning the abstract parity function for large numbers exacerbates an already fatal condition. The network’s inability to generalize to unseen positions adds profound noise to the learning process, which is already struggling to handle the non-associative nature of the game, creating a vicious cycle that makes true expert-level mastery impossible.
@@
    The Champion-Expert Dichotomy as a New Evaluation Metric: Our work highlights that winning a game and truly understanding it are not the same. We propose that the distinction between a "champion" (an agent that wins from the start state) and an "expert" (an agent that plays optimally from all states) is a crucial and under-explored metric for evaluating the true mastery and generalization capabilities of reinforcement learning algorithms. This distinction is especially critical for domains where adversarial attacks or rare, but strategically important, positions can expose the limitations of an agent’s learned knowledge.
@@ Acknowledgements
This work was funded by the Chinese Scholarship Council (CSC). We appreciate the assistance of the IT Research team at Queen Mary University of London for supporting us in using the Apocrita HPC facility. We would also like to thank the LCZero development team for providing detailed instructions for the use and explaining the inner workings of LCZero. We are also grateful for the comments we received from Dr. Ankur Mali and other researchers, which helped us greatly improve our paper.
Declarations
The authors declare that they have no conflict of interest.

### v6

@@ 6 Concluding remarks and conjectures
From a cognitive science perspective, the difference between games like Chess (or Go) and Nim is not merely one of complexity but of the very nature of optimal strategy. Mastering Chess and Go primarily involves associative pattern matching, heuristic evaluation of board configurations, and deep subsymbolic calculation. A human Grandmaster’s mastery is built upon millions of learned patterns, making it an ideal domain for reinforcement learning agents that excel at function approximation on large, but structured, data distributions. In contrast, Nim is governed by a single, elegant mathematical principle: the nim-sum. A human’s mastery of Nim is non-incremental; it is achieved not through extensive pattern recognition but through a single, abstract insight into combinatorial game theory. This disparity highlights a crucial distinction: AlphaZero excels at learning through association, but fails when a problem requires a form of symbolic reasoning that cannot be implicitly learned from the correlation between game states and outcomes.
@@
Our work empirically demonstrates that this theoretical vulnerability becomes a tangible, catastrophic failure mode in practice. While the problem of a neural network’s difficulty in learning parity from a uniform distribution is a well-established theoretical finding, our contribution is to show how this representational bottleneck manifests and breaks the AlphaZero learning loop in a real-world, self-play setting. The motivation behind our research is not to simply identify a known weakness, but to meticulously dissect its consequences and chart a path for extending the AlphaZero paradigm. By analysing the breakdown of the policy and value networks on specific Nim positions, we have revealed how a critical representational gap can trigger a cascade of failures.
@@
The core of the problem, as our analysis shows, is a breakdown in the crucial positive feedback loop that underpins reinforcement learning. The agent’s performance on small boards, while seemingly successful, is based on a precarious "champion" notion of mastery, memorising a few optimal opening moves. This allows the agent to steer the game into a sub-space it has seen frequently. However, as the number of heaps increases to just seven, the policy network’s predictions become near-uniform, and the value network’s assessments drift to a coin-toss probability. This lack of a strong initial policy prevents the MCTS from exploring the state space effectively, leaving it unable to find winning moves, even with millions of simulations. This is not a failure of calculation; it is a failure of learned guidance. Crucially, this happens even on boards where the state space is still small enough that many states would be revisited during training. In this scenario, the learning process becomes completely derailed, as the agent fails to generalise to any new positions, regardless of whether the parity function is complex. When the state space becomes truly vast (e.g., with 50+ heaps), the difficulty of learning the abstract parity function for large numbers exacerbates an already fatal condition. The network’s inability to generalise to unseen positions adds profound noise to the learning process, which is already struggling to handle the non-associative nature of the game, creating a vicious cycle that makes true expert-level mastery impossible.
@@
    The Champion-Expert Dichotomy as a New Evaluation Metric: Our work highlights that winning a game and truly understanding it are not the same. We propose that the distinction between a "champion" (an agent that wins from the start state) and an "expert" (an agent that plays optimally from all states) is a crucial and under-explored metric for evaluating the true mastery and generalisation capabilities of reinforcement learning algorithms. This distinction is especially critical for domains where adversarial attacks or rare, but strategically important, positions can expose the limitations of an agent’s learned knowledge.
@@ Acknowledgements
We appreciate the assistance of the IT Research team at Queen Mary University of London for supporting us in using the Apocrita HPC facility. We would also like to thank the LCZero development team for providing detailed instructions for the use and explaining the inner workings of LCZero.